OpenStudy (anonymous):

It takes 7hrs for 5 cooks to prepare dinner. How long will it take 6 cooks to prepare a dinner?

OpenStudy (saifoo.khan):

too many cooks spoil the Broth!

OpenStudy (anonymous):

7/5 = x/6 cross multiply for 5x = 42 x=8 and 2/5 so 8 hours and 24 minutes but this problem doesnt really make much mathematic sense since cook time would be the same...

OpenStudy (anonymous):

Its actually 5.8 hrs

OpenStudy (anonymous):

how did you get that???

OpenStudy (anonymous):

Thats what my program said was the correct answer

OpenStudy (anonymous):

well your program needs to show work!

OpenStudy (radar):

First we can assume that it will take less than 7 hrs, can't we?

OpenStudy (anonymous):

If it takes 7 hours for 5 cooks to cook dinner, it takes 35 man-hours to cook a diner. 35 man-hours over 6 cooks is 35/6 hours to cook diner using 6 cooks.

OpenStudy (radar):

Thanks robtobey, that would work out to be about 5.8222 hours, which indeed is less than 7.

OpenStudy (radar):

I mean 5.8333 hrs

OpenStudy (radar):

Which so happens to agree with Needzhelp77.

OpenStudy (anonymous):

@ radar I believe you are one of the very early problem solvers on this site. Wasn't it great in the old day when there were only 1500 students?

OpenStudy (radar):

Yes, it seems to be a furious pace (just to keep up).

OpenStudy (radar):

Geez, I just saw the number, 22,386 students!

OpenStudy (anonymous):

I think today may be my last day here for a variety of reasons. Can't deal with the text talk and ambiguous problem statements. Very difficult to get medals now because there is some one hovering over a key board, transfixed on a computer screen waiting for the next problem to be posted. At least I think I got one problem right that no one even attempted. Refer to: http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e4a90dc0b8b9588048f0086

OpenStudy (radar):

Had to go and look at that link, I see no further comment, so it is more than likely correct. It is above my pay grade. I am reviewing these problems to sort of refresh my memory. I am retired now and I don't have to do any math anymore, but still like to dabble if you know what I mean.

OpenStudy (anonymous):

Same here. By the way, I may be one of the oldest seniors here. Turned 77 on aug 10.

OpenStudy (radar):

Sorry to see you go, as I recall you have been very helpful and guided many students. Concur with the hovering, sometimes I think the award system can be harmful. I would rather take time with one student than provide a 100 students with just an answer.

OpenStudy (anonymous):

You might pass on your thoughts to the site management via the openstudy feedback. May check in on occasion. Adios.

OpenStudy (radar):

Adios, sayanara, aufweiderzine,